OpenStudy (anonymous):

10.0 Points If you roll one dice and it lands on a 6, what is the percent chance of you rolling a 6 on another roll?

OpenStudy (anonymous):

i guess is is \(\frac{1}{6}\) or whatever that is as a percent

OpenStudy (anonymous):

IS IT 16%

OpenStudy (mathstudent55):

Each roll of a fair die is an independent event. The fact that you just rolled a 6 does not influence the outcome of the next roll. The probaility of rolling a 6 is still: P(Rolling 6) = 1/6 = 16.666...%

OpenStudy (anonymous):

okay thankuu very much